# Polynomial-time Extraction of Entanglement Resources

### v2

@@ I Introduction
A representative example is the Bell-VM problem, which is known to be NP-complete [6]. Bell-VM asks whether a given graph state can be transformed into a set of $k$ Bell pairs between specific node pairs, by using only local Clifford operations, Pauli measurements, and classical communication (LC + LPM + CC). Research in this area has potential applications in point-to-point quantum communication protocols, where the extracted $k$ Bell pairs can be used for the parallel transmission of $k$ informational qubits via teleporting protocol [2]. However, the locations of the $k$ node pairs must be chosen in advance, limiting the adaptability to the dynamics of the communication requests.
A related line of research focuses on determining whether a GHZ state can be extracted from a given graph state, by using only LC + LPM + CC operations. This problem has also been proven to be NP-complete [5]. For ease of reference, we refer to the aforementioned problem as GHZ-VM problem. Although the GHZ-VM problem concerns the extraction of a single GHZ state, it offers a significant advantage over the Bell-VM formulation, since it inherently supports adaptability. Indeed, a GHZ state allows the dynamic extraction of an EPR pair between any pair of nodes sharing the original GHZ state. And, notably, this extraction can occur at runtime, driven by the actual and potentially time-varying communication needs of the nodes.
@@
  • We propose a polynomial-time algorithm for the Remote-VM (see Sec. III-A), by exploiting graph-theory tools and only LC + LPM + CC.
@@ II-A Preliminaries
Given a $N$-qubit graph state $\left|G\right\rangle$ and its corresponding graph $G=(V,E)$, two network nodes $i$ and $j$ are defined as remote if the corresponding vertices $v_{i},v_{j}$ are non-adjacent in $G$, i.e., if:
@@
The remote $n$-Gability of an $N$-qubit graph state $\left|G\right\rangle$ quantifies the number of $n$-qubit GHZ states ($n\leq N$) that can be extracted among remote nodes using LOCC. The number of such GHZ states extractable from $\left|G\right\rangle$ is denoted by $r_{g}(n)$.
@@
The remote Pairability of an $N$-qubit graph state $\left|G\right\rangle$ quantifies the number of EPR pairs that can be extracted among remote nodes using LOCC. The number of extractable EPR pairs from a graph state $\left|G\right\rangle$ is denoted by $r_{g}(2)$.
@@
The term “opposite” in Def. 5 is used to highlight that the remote nodes belong to different partitions. In Fig. 3, we provide a pictorial representation for opposite remote sets. Clearly, vertices belonging to the same partition are remote per se, as a consequence of the two-colorable graph state definition. The opposite remote set can be extended from vertices to set of vertices belonging to the same partition. In such a case, by considering a subset of vertices $A\subseteq P_{i}$ we formally define opposite remote set $\overline{N}_{\cup}(A)$ as the union of opposite remote set of each vertex in $A$, and $\overline{N}_{\cap}(A)$ as the intersection among the opposite remote sets of each vertex in $A$, i.e.:
@@ III-B Algorithm Complexity Analysis
Remarkably, our algorithm leverages graph theory and uses only single-qubit Clifford operations, single-qubit Pauli measurement and classical communication (LC + LPM + CC). Theorem 1 shows that our algorithm can quickly determine the extractable volume. Furthermore, the final volume, obtained through successive updates, represents an enhanced extractable value, as shown in (9). We refer interested readers to Appendix B for the proof.
@@ IV Performance Evaluation
Figure 6: Pictorial illustration for Remote extraction from general graph state. (1) The sample Protein-Protein Interactions (PPI) topology is generated by NetworkX library. (2) The extract bipartite subgraphs from PPI topology with fixed number of nodes. (3) The remote extracted entanglement resource, i.e., a $4$-qubit GHZ and a $3$-qubit GHZ, can be obtained by Algorithm 1.
@@ IV-B1 Setup
We evaluate the extractable values against different bipartite graph structures by randomly varying the number of edges $m$, while keeping the total number of nodes constant and equal to $50$. This allows for a fair comparison across various graph instances. Furthermore, for the sake of generality, we distribute the nodes in different ways: one approach allocates nodes unequally across partitions, while the other ensures an equal number of nodes in each partition. More into details, we consider graphs with partitions ($P_{1},P_{2}$) having sizes (10, 40), (20, 30), (25, 25), respectively.

### v3

@@ I Introduction
A representative example is the Bell-VM problem, which is known to be NP-complete [6]. Bell-VM asks whether a given graph state can be transformed into a set of $k$ Bell pairs between specific node pairs, by using only local Clifford operations, Pauli measurements, and classical communication (LC + LPM + $C$). Research in this area has potential applications in point-to-point quantum communication protocols, where the extracted $k$ Bell pairs can be used for the parallel transmission of $k$ informational qubits via teleporting protocol [2]. However, the locations of the $k$ node pairs must be chosen in advance, limiting the adaptability to the dynamics of the communication requests.
A related line of research focuses on determining whether a GHZ state can be extracted from a given graph state, by using only LC + LPM + $C$ operations. This problem has also been proven to be NP-complete [5]. For ease of reference, we refer to the aforementioned problem as GHZ-VM problem. Although the GHZ-VM problem concerns the extraction of a single GHZ state, it offers a significant advantage over the Bell-VM formulation, since it inherently supports adaptability. Indeed, a GHZ state allows the dynamic extraction of an EPR pair between any pair of nodes sharing the original GHZ state. And, notably, this extraction can occur at runtime, driven by the actual and potentially time-varying communication needs of the nodes.
@@
  • We propose a polynomial-time algorithm for the Remote-VM (see Sec. III-A), by exploiting graph-theory tools and only LC + LPM + $C$.
@@ II-A Preliminaries
Given a NN-qubit graph state $\left|G\right\rangle$ and its corresponding graph $G=(V,E)$, two network nodes $i$ and $j$ are defined as remote if the corresponding vertices $v_{i},v_{j}$ are non-adjacent in $G$, i.e., if:
@@
The remote $n$-Gability of an NN-qubit graph state $\left|G\right\rangle$ quantifies the number of $n$-qubit GHZ states ($n\leq N$) that can be extracted among remote nodes using LOCC. The number of such GHZ states extractable from $\left|G\right\rangle$ is denoted by $r_{g}(n)$.
@@
The remote Pairability of an NN-qubit graph state $\left|G\right\rangle$ quantifies the number of EPR pairs that can be extracted among remote nodes using LOCC. The number of extractable EPR pairs from a graph state $\left|G\right\rangle$ is denoted by $r_{g}(2)$.
@@
The term “opposite” in Def. 5 is used to highlight that the remote nodes belong to different partitions. In Fig. 3, we provide a pictorial representation for opposite remote sets. Clearly, vertices belonging to the same partition are remote per se, as a consequence of the two-colorable graph state definition. The opposite remote set can be extended from vertices to set of vertices belonging to the same partition. In such a case, by considering a subset of vertices $A\subseteq P_{i}$ we formally define opposite remote set $\overline{N}_{\cup}(A)$ as the union of opposite remote set of each vertex in AA, and $\overline{N}_{\cap}(A)$ as the intersection among the opposite remote sets of each vertex in AA, i.e.:
@@ III-B Algorithm Complexity Analysis
Remarkably, our algorithm leverages graph theory and uses only single-qubit Clifford operations, single-qubit Pauli measurement and classical communication (LC + LPM + $C$). Theorem 1 shows that our algorithm can quickly determine the extractable volume. Furthermore, the final volume, obtained through successive updates, represents an enhanced extractable value, as shown in (9). We refer interested readers to Appendix B for the proof.
@@ IV Performance Evaluation
Figure 6: Pictorial illustration for Remote extraction from general graph state. (1) The sample Protein-Protein Interactions (PPI) topology is generated by NetworkX library. (2) The extract bipartite subgraphs from PPI topology with fixed number of nodes. (3) The remote extracted entanglement resource, i.e., a 44-qubit GHZ and a 33-qubit GHZ, can be obtained by Algorithm 1.
@@ IV-B1 Setup
We evaluate the extractable values against different bipartite graph structures by randomly varying the number of edges $m$, while keeping the total number of nodes constant and equal to 5050. This allows for a fair comparison across various graph instances. Furthermore, for the sake of generality, we distribute the nodes in different ways: one approach allocates nodes unequally across partitions, while the other ensures an equal number of nodes in each partition. More into details, we consider graphs with partitions ($P_{1},P_{2}$) having sizes (10, 40), (20, 30), (25, 25), respectively.